# Aligning Text to Image in Diffusion Models is Easier Than You Think

### v4

@@ Appendix E Additional Results on T2I Generation
Additional qualitative results for SD1.5 and SDXL on the COCO-val5K dataset are presented in fig. 7 and fig. 8, respectively. Furthermore, qualitative examples on the GenEval benchmark are provided in fig. 9.
@@ Appendix F Additional Results on Image Editing
Qualitative comparisons of SoftREPA across various editing methods, including PnP [37], MasaCtrl [4] with DDIM [34], Direct Inversion [14], and FlowEdit [19], on the PIEBench dataset are presented in fig. 11, corresponding to the quantitative results in table 2.
Quantitative results on the DIV2K and Cat2Dog datasets under varying target CFG scales are reported in table 5 and table 6, respectively. The corresponding qualitative results for different CFG scales are also shown in table 5. Additionally, qualitative examples across different numbers of editing steps are provided in fig. 12 and fig. 10.
@@
Figure 12: The additional qualitative results on editing using FlowEdit. The edited1-3 using soft tokens are vary on the number of editing steps, $27,29,30$ out of $50$ total timesteps, respectively.
@@ Appendix G Additional Results on Ablation Study
Quantitative results for varying the number of layers and soft tokens are shown in table 7, with corresponding qualitative examples in fig. 13 and fig. 14. As seen in fig. 13, text prompts are increasingly emphasized as more layers adopt soft tokens. The same input image is used in fig. 14 to illustrate the effect of different token counts.

### v5

@@ Appendix E Additional Results on T2I Generation
Additional qualitative results for SD1.5 and SDXL on the COCO-val5K dataset are presented in fig.˜7 and fig.˜8, respectively. Furthermore, qualitative examples on the GenEval benchmark are provided in fig.˜9.
@@ Appendix F Additional Results on Image Editing
Qualitative comparisons of SoftREPA across various editing methods, including PnP [38], MasaCtrl [4] with DDIM [35], Direct Inversion [14], and FlowEdit [19], on the PIEBench dataset are presented in fig.˜11, corresponding to the quantitative results in table˜2.
Quantitative results on the DIV2K and Cat2Dog datasets under varying target CFG scales are reported in table˜8 and table˜9, respectively. The corresponding qualitative results for different CFG scales are also shown in table˜8. Additionally, qualitative examples across different numbers of editing steps are provided in fig.˜12 and fig.˜10.
@@
Figure 12: The additional qualitative results on editing using FlowEdit. The edited1-3 using soft tokens are vary on the number of editing steps, $27,29,30$ out of 5050 total timesteps, respectively.
@@ Appendix G Additional Results on Ablation Study
Quantitative results for varying the number of layers and soft tokens are shown in table˜10, with corresponding qualitative examples in fig.˜13 and fig.˜14. As seen in fig.˜13, text prompts are increasingly emphasized as more layers adopt soft tokens. The same input image is used in fig.˜14 to illustrate the effect of different token counts.
@@
Appendix H Additional Results on comparing with LoRA
In table˜11, we conducted various conditions on LoRA finetuning, including top-5-layer and full-layer LoRA, as well as the corresponding number of trainable parameters and the chosen LoRA rank. Specifically, LoRA with top-5 layers uses 0.4M parameters, LoRA applied to all layers uses 2.3M parameters, and our soft token approach uses 0.9M parameters for 5 layers. We adopt a LoRA rank of 4, which is commonly used for diffusion model finetuning.
To enable a direct comparison with a similar number of parameters, we additionally report results for LoRA applied to the top-10 layers (0.9M parameters) and top-5 layers with rank 8 (0.9M parameters). Among these settings, LoRA on the top-5 layers with rank 4 yields the best results within LoRA baselines. Importantly, the combination of Soft Tokens and contrastive Loss consistently achieves the strongest performance across key metrics, supporting our claim that both techniques independently and meaningfully improve text-image alignment.
Human Preference Text Alignment Image Quality
[…]
SD3 90.89 22.50 26.26 27.97 56.87 42.47
SD3 + LoRA + Contrastive loss 4 1-5 0.4M 97.72 22.57 26.42 28.44 70.33 42.71
SD3 + LoRA + Contrastive loss 8 1-5 0.9M 95.92 22.55 26.28 28.28 70.30 42.55
SD3 + LoRA + Contrastive loss 4 1-10 0.9M 92.72 22.54 26.23 28.10 70.53 42.59
SD3 + LoRA + Contrastive loss 4 all layers 2.3M 80.89 22.20 26.70 25.97 71.57 41.54
SD3 + Soft tokens + Contrastive loss (Ours) - 1-5 0.9M 106.32 22.49 26.92 28.70 60.76 42.99
Table 11: Quantitative comparison of T2I generation between soft tokens and various configurations of LoRA on SD3. Generation quality is evaluated on the COCO-val 1K [23]. ImageReward, CLIP, HPS, and LPIPS are scaled by $\times 10^{2}$.